Activity 3: Gravitational Force



The mass of m1 of one of the small sheres of a Cavendish balance is 0.02 kg, the



mass m2 of the nearest large sphere is 0.065 kg, and the center-to-center distance



between them is 0.04 m. Find the gravitational force Fg on each sphere due to the other.

1
Expert's answer
2021-11-23T17:18:42-0500

"m_1=0.02kg\\m_2=0.065kg\\r=0.04m"

Gravitation force

"F=\frac{Gm_1m_2}{r^2}"


"F=\frac{6.67\times10^{-11}\times0.02\times0.065}{0.04^2}=5.41\times10^{-11}N"